Junction Rule:

I1+I2=I3                     (1)

Loop Rule: Loop ABEFA

ε1-I1R1-I3R3=0          2

Loop Rule: Loop CDEFC

ε2-I2R2-I3R3=0          (3)

HOW DO YOU FIND THE CALCULATED CURRENT PLEASE SHOW THE STEPS! THANK YOUData Table 2 1 2 3 4 5 Measured Voltage Calculated Current (Using Ohms Law) Measured Current Calculated Current (Using Equat
Answer #1


U I + I = I3 - 0 2.- I, R. - I2R3 = 0 -> I = {. - Iz R3 R celso E - I R₂ - Iz R₂ = 0 I = E2-J3R3 > B putting I & I in S I R₂